OpenStudy (curry):

Find all zeros of f(x) = x^4 - 6x^3+5x^2 + 22x-40 given (2-i) is a zero

OpenStudy (phi):

if 2-i is a zero then you know 2+i is also a root. complex roots come in complex conjugate pairs. so you could multiply out (x-(2-i))(x-(2+i)) to get a quadratic and divide that into the original expression to get another quadratic. factor this second quadratic for the remaining 2 zeros

OpenStudy (curry):

ye but i keep getting a error in my calculations

OpenStudy (anonymous):

you could multiply out (x-(2-i))(x-(2+i)) to get a quadratic = x^2-4x+5

OpenStudy (curry):

ye and then itrs the factoreing to 6th degreee that messes me up

OpenStudy (phi):

you divide x^2-4x+5 into the quartic. It will go in evenly, a give you a quadratic that can be factored.

OpenStudy (anonymous):

divide: x^2-4x+5 | x^4 - 6x^3+5x^2 + 22x-40 goes in on the first time through x^2 times...

OpenStudy (anonymous):

X^2 times x^2-4x+5 =... then subtract, like in long division...

OpenStudy (anonymous):

\[\frac{x^4-6 x^3+5 x^2+22 x-40}{x^2-4 x+5}=x^2-2 x-8 \]